# Dusit Niyato, , Bo Ai

### v1

@@ IV Optimal Bilinear Equalizer Design
The combining schemes in (7) and (13) can be named as BE. This is because the decoding estimates in (4) and (11) are not only linear to the received data vectors $\mathbf{y}_{m}$ and $\mathbf{y}_{mk}$ but also linear to the channel estimates $\hat{\mathbf{g}}_{k}$ and $\hat{\mathbf{g}}_{mk}$, for the centralized and distributed processing schemes, respectively [24]. In this section, we will optimally design $\mathbf{W}_{k}$ and $\mathbf{W}_{mk}$ to derive the OBE which can achieve the best achievable SE performance.
@@ IV-A C-OBE Combining Scheme Design
For the centralized processing scheme, we optimally design the C-OBE, which can maximize the achievable SE for the centralized processing scheme based on the UatF capacity bound in Sec. III-A. When applying the BE combining scheme (7), we can represent the UatF bound based SINR in (6) as
$\mathrm{SINR}_{k}^{\mathrm{c},\mathrm{UatF}}=$ (22)
$\frac{p_{k}|\mathbb{E}\{\hat{\mathbf{g}}_{k}^{H}\mathbf{W}_{k}^{H}\mathbf{g}_{k}\}|^{2}}{\sum\limits_{l=1}^{K}{p_{l}\mathbb{E}\{|\hat{\mathbf{g}}_{k}^{H}\mathbf{W}_{k}^{H}\mathbf{g}_{l}|^{2}\}}\!\!-\!\!p_{k}|\mathbb{E}\{\hat{\mathbf{g}}_{k}^{H}\mathbf{W}_{k}^{H}\mathbf{g}_{k}\}|^{2}\!+\!\sigma^{2}\mathbb{E}\{\parallel\mathbf{W}_{k}\hat{\mathbf{g}}_{k}\parallel^{2}\}}.$
Furthermore, we can derive the optimal $\mathbf{W}_{k}$, which can maximize (6) as follows [^2] In the following, we focus on the analysis of the SINR. The achievable SE can be easily derived from the SINR based on the above corollaries. Notably, when the SINR is maximized, the achievable SE can also be maximized since $\tau_{c}$ and $\tau_{p}$ are constants in this paper..
@@
$\mathrm{SINR}_{k}^{\mathrm{c},\mathrm{UatF},*}=p_{k}\mathrm{vec}(\mathbb{E}\{\mathbf{g}_{k}\hat{\mathbf{g}}_{k}^{H}\})^{H}\mathbf{w}_{k}^{*},$ (23)
@@
$\overline{\mathbf{w}}_{k}^{*}=\mathbf{\Gamma}_{k}^{-1}\bar{\mathbf{r}}_{k},$ (25)
where $\mathbf{\Gamma}_{k}=\sum\nolimits_{l=1}^{K}{p_{l}[(\bar{\mathbf{G}}_{kk}^{T}\otimes\bar{\mathbf{G}}_{ll})+(\bar{\mathbf{G}}_{kk}^{T}\otimes\check{\mathbf{R}}_{l})}+(\hat{\mathbf{R}}_{k}^{T}\otimes\bar{\mathbf{G}}_{ll})+(\hat{\mathbf{R}}_{k}^{T}\otimes\check{\mathbf{R}}_{l})]+\sum\nolimits_{l\in\mathcal{P}_{k}}{p_{k}p_{l}^{2}\tau_{p}^{2}\tilde{\mathbf{r}}_{lk}\tilde{\mathbf{r}}_{lk}^{H}}+p_{k}^{2}\tau_{p}\bar{\mathbf{g}}_{kk}\tilde{\mathbf{r}}_{kk}^{H}+p_{k}^{2}\tau_{p}\tilde{\mathbf{r}}_{kk}\bar{\mathbf{g}}_{kk}^{H}+p_{k}\bar{\mathbf{\Upsilon}}_{k}-p_{k}(\bar{\mathbf{G}}_{kk}^{T}\otimes\bar{\mathbf{G}}_{kk})-p_{k}\bar{\mathbf{r}}_{k}\bar{\mathbf{r}}_{k}^{H}+\sigma^{2}(\bar{\mathbf{R}}_{k}^{T}\otimes\mathbf{I}_{MN})\in\mathbb{C}^{M^{2}N^{2}\times M^{2}N^{2}}$ with $\tilde{\mathbf{r}}_{lk}=\mathrm{vec}(\check{\mathbf{R}}_{l}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{k})$, $\bar{\mathbf{g}}_{kk}=\mathrm{vec}(\bar{\mathbf{G}}_{kk})$, and $\bar{\mathbf{r}}_{k}=\mathrm{vec}(\bar{\mathbf{R}}_{k})$. As for $\bar{\mathbf{\Upsilon}}_{k}\in\mathbb{C}^{M^{2}N^{2}\times M^{2}N^{2}}$, the $(a,b)$-th element of the $(c,d)$-th block matrix of it $\bar{\mathbf{\Upsilon}}_{k,cd}\in\mathbb{C}^{MN\times MN}$ with $\{a,b,c,d\}=1,\dots,MN$ is $[\bar{\mathbf{\Upsilon}}_{k,cd}]_{ab}=\bar{g}_{m_{1}k,n_{1}}\bar{g}_{m_{2}k,n_{2}}^{*}\bar{g}_{m_{3}k,n_{3}}\bar{g}_{m_{4}k,n_{4}}^{*}$ if $m_{1}=m_{2}$, $m_{3}=m_{4}$ or $m_{1}=m_{4}$, $m_{2}=m_{3}$, where $\{m_{1},m_{2},m_{3},m_{4}\}=1,\dots,M$, $\{n_{1},n_{2},n_{3},n_{4}\}=1,\dots,N$, $d=(m_{1}-1)N+n_{1}$, $c=(m_{2}-1)N+n_{2}$, $a=(m_{3}-1)N+n_{3}$, $b=(m_{4}-1)N+n_{4}$, and $\bar{g}_{m_{1}k,n_{1}}$ is the $n_{1}$-th element of $\bar{\mathbf{g}}_{m_{1}k}$, respectively. The maximum value of $\mathrm{SINR}_{k}^{\mathrm{c},\mathrm{UatF},*}$ can also be computed in closed-form as $\overline{\mathrm{SINR}}_{k}^{\mathrm{c},\mathrm{UatF},*}=p_{k}\bar{\mathbf{r}}_{k}^{H}\overline{\mathbf{w}}_{k}^{*}$.
@@
Note that the considered random phase-shifts considered result in the phenomenon that some cross terms from different APs or different UEs become 0. When the phase-shifts are assumed to be perfectly mitigated through all processing phases in the system, so the channel model in Sec. II-A reduces to $\mathbf{g}_{mk}=\bar{\mathbf{g}}_{mk}+\check{\mathbf{g}}_{mk}$, the closed-form results in Theorem 1 and Theorem 3 would have different forms. More specifically, we obtain the expressions in Theorem 1, but using $\bar{\mathbf{G}}_{kk}=\bar{\mathbf{g}}_{k}^{\prime}\bar{\mathbf{g}}_{k}^{\prime,H}$ with $\bar{\mathbf{g}}_{k}^{\prime}\triangleq[\bar{\mathbf{g}}_{1k}^{T},\dots,\bar{\mathbf{g}}_{Mk}^{T}]^{T}$,
$\varepsilon_{kl}$ $=\sqrt{p_{k}p_{l}}\tau_{p}\mathrm{tr}(\mathbf{W}_{k}^{H}\bar{\mathbf{G}}_{lk})\mathrm{tr}(\mathbf{W}_{k}\check{\mathbf{R}}_{k}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{l})$
$+\sqrt{p_{k}p_{l}}\tau_{p}\mathrm{tr}(\mathbf{W}_{k}\bar{\mathbf{G}}_{kl})\mathrm{tr}(\mathbf{W}_{k}^{H}\check{\mathbf{R}}_{l}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{k})$
@@
with $\bar{\mathbf{G}}_{lk}=\bar{\mathbf{g}}_{l}^{\prime}\bar{\mathbf{g}}_{k}^{\prime,H}$, and $\omega_{k}=0$, respectively. Furthermore, the expressions in Theorem 3 hold but using $\mathbf{\Gamma}_{k}=\sum_{l=1}^{K}{p_{l}[(\bar{\mathbf{G}}_{kk}^{T}\otimes\bar{\mathbf{G}}_{ll})+(\bar{\mathbf{G}}_{kk}^{T}\otimes\check{\mathbf{R}}_{l})+(\hat{\mathbf{R}}_{k}^{T}\otimes\bar{\mathbf{G}}_{ll})}+(\hat{\mathbf{R}}_{k}^{T}\otimes\check{\mathbf{R}}_{l})]+\sum_{l\in\mathcal{P}_{k}}{p_{l}[\sqrt{p_{k}p_{l}}\tau_{p}\bar{\mathbf{g}}_{lk}\tilde{\mathbf{r}}_{lk}^{H}+\sqrt{p_{k}p_{l}}\tau_{p}\tilde{\mathbf{r}}_{lk}\bar{\mathbf{g}}_{lk}^{H}}+p_{k}p_{l}\tau_{p}^{2}\tilde{\mathbf{r}}_{lk}\tilde{\mathbf{r}}_{lk}^{H}]-p_{k}\bar{\mathbf{r}}_{k}\bar{\mathbf{r}}_{k}^{H}+\sigma^{2}(\bar{\mathbf{R}}_{k}^{T}\otimes\mathbf{I}_{MN})$, where $\bar{\mathbf{g}}_{lk}=\mathrm{vec}(\bar{\mathbf{G}}_{lk})$. These results can be easily derived following the same methods as in Appendix B and Appendix E. As observed, if the phase-shifts are perfectly mitigated, some cross terms from different APs or different UEs become non-zero, such as $\bar{\mathbf{G}}_{lk}$. Note that the results for the distributed processing scheme, including the achievable SE, DG-OBE, and DL-OBE expressions in scenarios with perfectly mitigated phase-shifts, can also be derived using the same approach as this remark. Therefore, they are omitted in the subsequent discussion for brevity.
@@ IV-B DG-OBE Combining Scheme Design
In this subsection, we study the DG-OBE combining scheme design to maximize (14). We define $\mathbf{w}_{k}^{\mathrm{d}}=[\mathbf{w}_{1k}^{T},\mathbf{w}_{2k}^{T},\dots,\mathbf{w}_{Mk}^{T}]^{T}\in\mathbb{C}^{MN^{2}}$, where $\mathbf{w}_{mk}=\mathrm{vec}(\mathbf{W}_{mk})\in\mathbb{C}^{N^{2}}$. To separate the optimization of $\{\mathbf{W}_{mk}:m=1,\dots,M\}$ and $\mathbf{a}_{k}$, we first assume all elements in $\mathbf{a}_{k}$ to be equally and thus, $\mathrm{SINR}_{k}^{\mathrm{LSFD}}$ in (15) reduces to (26) on the top of the previous page. Based on (26), we can optimize the distributed BE matrices $\{\mathbf{W}_{mk}:m=1,\dots,M\}$ to maximize the achievable SE performance for UE $k$ in (26) as follows.
@@
The optimal matrices $\{\mathbf{W}_{mk}^{*}:m=1,\dots,M\}$, maximizing the effective SINR for the distributed processing scheme as in (15), can be computed as $\mathbf{W}_{mk}^{*}=\mathrm{vec}^{-1}(\mathbf{w}_{mk}^{*})\in\mathbb{C}^{N\times N}$ with $\mathbf{w}_{mk}^{\mathrm{d},*}=[\mathbf{w}_{k}^{\mathrm{d},*}]_{(m-1)N^{2}+1:mN^{2}}\in\mathbb{C}^{N^{2}}$ and
$\mathbf{w}_{k}^{\mathrm{d},*}\!\!\!=\!\!\left(\sum_{l=1}^{K}{p_{l}\mathbb{E}\{\mathbf{z}_{kl}\mathbf{z}_{kl}^{H}\}\!-\!p_{k}\mathbb{E}\{\mathbf{h}_{k}\}\mathbb{E}\{\mathbf{h}_{k}\}^{H}\!+\!\sigma^{2}\mathbf{\Theta}_{k}}\!\!\right)^{-1}\!\!\!\mathbb{E}\{\mathbf{h}_{k}\},$ (27)
where $\mathbf{z}_{kl}=[\mathrm{vec}(\mathbf{g}_{1l}\hat{\mathbf{g}}_{1k}^{H})^{T},\dots,\mathrm{vec}(\mathbf{g}_{Ml}\hat{\mathbf{g}}_{Mk}^{H})^{T}]^{T}\in\mathbb{C}^{MN^{2}}$, $\mathbf{h}_{k}=[\mathrm{vec}(\mathbf{g}_{1k}\hat{\mathbf{g}}_{1k}^{H})^{T},\dots,\mathrm{vec}(\mathbf{g}_{Mk}\hat{\mathbf{g}}_{Mk}^{H})^{T}]^{T}\in\mathbb{C}^{MN^{2}}$, and $\mathbf{\Theta}_{k}=\mathrm{diag}\{(\mathbb{E}\{\hat{\mathbf{g}}_{1k}\hat{\mathbf{g}}_{1k}^{H}\}^{T}\otimes\mathbf{I}_{N}),\dots,(\mathbb{E}\{\hat{\mathbf{g}}_{Mk}\hat{\mathbf{g}}_{Mk}^{H}\}^{T}\otimes\mathbf{I}_{N})\}\in\mathbb{C}^{MN^{2}\times MN^{2}}$. When $\{\mathbf{W}_{mk}^{*}:m=1,\dots,M\}$ are applied, the SINR in (26) can achieve the maximum value as $\widetilde{\mathrm{SINR}}_{k}^{\mathrm{LSFD},*}=p_{k}\mathbb{E}\{\mathbf{h}_{k}\}^{H}\mathbf{w}_{k}^{\mathrm{d},*}$.
@@
The closed-form results of $\mathbf{w}_{k}^{\mathrm{d},*}$ can be computed as $\overline{\mathbf{w}}_{k}^{\mathrm{d},*}=\mathbf{\Lambda}_{k}^{-1}\overline{\mathbf{h}}_{k}$, where $\mathbf{\Lambda}_{k}=\sum_{l=1}^{K}{\mathbf{\Lambda}_{kl}^{(1)}}+\sum_{l\in\mathcal{P}_{k}}{(\mathbf{\Lambda}_{kl}^{(2)}+\mathbf{\Lambda}_{kl}^{(3)})}-p_{k}\overline{\mathbf{h}}_{k}\overline{\mathbf{h}}_{k}^{H}+\sigma^{2}\mathbf{\Lambda}_{k}^{(4)}\in\mathbb{C}^{MN^{2}\times MN^{2}}$ and $\overline{\mathbf{h}}_{k}=[\overline{\mathbf{h}}_{1k}^{T},\dots,\overline{\mathbf{h}}_{Mk}^{T}]^{T}\in\mathbb{C}^{MN^{2}}$ with $\overline{\mathbf{h}}_{mk}=\mathrm{vec}(\bar{\mathbf{R}}_{mk})\in\mathbb{C}^{N^{2}}$. Besides, $\mathbf{\Lambda}_{kl}^{(1)}=\mathrm{diag}(\mathbf{\Lambda}_{1kl}^{(1)},\dots,\mathbf{\Lambda}_{Mkl}^{(1)})$, $\mathbf{\Lambda}_{kl}^{(2)}=\mathrm{diag}(\mathbf{\Lambda}_{1kl}^{(2)},\dots,\mathbf{\Lambda}_{Mkl}^{(2)})$, $\mathbf{\Lambda}_{k}^{(4)}=\mathrm{diag}(\mathbf{\Lambda}_{1k}^{(4)},\dots,\mathbf{\Lambda}_{Mk}^{(4)})$ are all diagonal block matrices with dimension $\mathbb{C}^{MN^{2}\times MN^{2}}$, where $\mathbf{\Lambda}_{mkl}^{(1)}=p_{l}[(\bar{\mathbf{G}}_{mkk}^{T}\otimes\bar{\mathbf{G}}_{mll})+(\bar{\mathbf{G}}_{mkk}^{T}\otimes\check{\mathbf{R}}_{ml})+(\hat{\mathbf{R}}_{mk}^{T}\otimes\bar{\mathbf{G}}_{mll})+(\hat{\mathbf{R}}_{mk}^{T}\otimes\check{\mathbf{R}}_{ml})]$,

### v2

@@ IV Optimal Bilinear Equalizer Design
The combining schemes in (7) and (13) can be named as BE. This is because the decoding estimates in (4) and (11) are not only linear to the received data vectors $\mathbf{y}_{m}$ and $\mathbf{y}_{mk}$ but also linear to the channel estimates $\widehat{\mathbf{g}}_{k}$ and $\widehat{\mathbf{g}}_{mk}$, for the centralized and distributed processing schemes, respectively [27]. In this section, we will optimally design $\mathbf{W}_{k}$ and $\mathbf{W}_{mk}$ to derive the OBE which can achieve the best achievable SE performance.
@@ IV-A C-OBE Combining Scheme Design
For the centralized processing scheme, we optimally design the C-OBE, which can maximize the achievable SE for the centralized processing scheme based on the UatF capacity bound in Sec. III-A. When applying the BE combining scheme (7), we can represent the UatF bound based SINR in (6) in (22) on the top of the previous page. Furthermore, we can derive the optimal $\mathbf{W}_{k}$, which can maximize (6) as follows [^2] In the following, we focus on the analysis of the SINR. The achievable SE can be easily derived from the SINR based on the above results. Notably, when the SINR is maximized, the achievable SE can also be maximized since $\tau_{c}$ and $\tau_{p}$ are constants in this paper..
@@
$\mathrm{SINR}_{k}^{\mathrm{c},\mathrm{UatF},*}=p_{k}\mathrm{vec}(\mathbb{E}\{\mathbf{g}_{k}\widehat{\mathbf{g}}_{k}^{H}\})^{H}\mathbf{w}_{k}^{*},$ (23)
@@
$\overline{\mathbf{w}}_{k}^{*}=\mathbf{\Gamma}_{k}^{-1}\overline{\mathbf{r}}_{k},$ (25)
where $\mathbf{\Gamma}_{k}=\sum\nolimits_{l=1}^{K}{p_{l}[(\overline{\mathbf{G}}_{kk}^{T}\otimes\overline{\mathbf{G}}_{ll})+(\overline{\mathbf{G}}_{kk}^{T}\otimes\check{\mathbf{R}}_{l})}+(\widehat{\mathbf{R}}_{k}^{T}\otimes\overline{\mathbf{G}}_{ll})+(\widehat{\mathbf{R}}_{k}^{T}\otimes\check{\mathbf{R}}_{l})]+\sum\nolimits_{l\in\mathcal{P}_{k}}{p_{k}p_{l}^{2}\tau_{p}^{2}\tilde{\mathbf{r}}_{lk}\tilde{\mathbf{r}}_{lk}^{H}}+p_{k}^{2}\tau_{p}\overline{\mathbf{g}}_{kk}\tilde{\mathbf{r}}_{kk}^{H}+p_{k}^{2}\tau_{p}\tilde{\mathbf{r}}_{kk}\overline{\mathbf{g}}_{kk}^{H}+p_{k}\overline{\mathbf{\Upsilon}}_{k}-p_{k}(\overline{\mathbf{G}}_{kk}^{T}\otimes\overline{\mathbf{G}}_{kk})-p_{k}\overline{\mathbf{r}}_{k}\overline{\mathbf{r}}_{k}^{H}+\sigma^{2}(\overline{\mathbf{R}}_{k}^{T}\otimes\mathbf{I}_{MN})\in\mathbb{C}^{M^{2}N^{2}\times M^{2}N^{2}}$ with $\tilde{\mathbf{r}}_{lk}=\mathrm{vec}(\check{\mathbf{R}}_{l}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{k})$, $\overline{\mathbf{g}}_{kk}=\mathrm{vec}(\overline{\mathbf{G}}_{kk})$, and $\overline{\mathbf{r}}_{k}=\mathrm{vec}(\overline{\mathbf{R}}_{k})$. As for $\overline{\mathbf{\Upsilon}}_{k}\in\mathbb{C}^{M^{2}N^{2}\times M^{2}N^{2}}$, the $(a,b)$-th element of the $(c,d)$-th block matrix of it $\overline{\mathbf{\Upsilon}}_{k,cd}\in\mathbb{C}^{MN\times MN}$ with $\{a,b,c,d\}=1,\dots,MN$ is $[\overline{\mathbf{\Upsilon}}_{k,cd}]_{ab}=\overline{g}_{m_{1}k,n_{1}}\overline{g}_{m_{2}k,n_{2}}^{*}\overline{g}_{m_{3}k,n_{3}}\overline{g}_{m_{4}k,n_{4}}^{*}$ if $m_{1}=m_{2}$, $m_{3}=m_{4}$ or $m_{1}=m_{4}$, $m_{2}=m_{3}$, where $\{m_{1},m_{2},m_{3},m_{4}\}=1,\dots,M$, $\{n_{1},n_{2},n_{3},n_{4}\}=1,\dots,N$, $d=(m_{1}-1)N+n_{1}$, $c=(m_{2}-1)N+n_{2}$, $a=(m_{3}-1)N+n_{3}$, $b=(m_{4}-1)N+n_{4}$, and $\overline{g}_{m_{1}k,n_{1}}$ is the $n_{1}$-th element of $\overline{\mathbf{g}}_{m_{1}k}$, respectively. $\mathrm{SINR}_{k}^{\mathrm{c},\mathrm{UatF},*}$ can also be computed in closed-form as $\overline{\mathrm{SINR}}_{k}^{\mathrm{c},\mathrm{UatF},*}=p_{k}\overline{\mathbf{r}}_{k}^{H}\overline{\mathbf{w}}_{k}^{*}$.
@@
Note that the considered random phase-shifts considered result in the phenomenon that some cross terms from different APs or different UEs become 0. When the phase-shifts are assumed to be perfectly mitigated through all processing phases in the system, so the channel model in Sec. II-A reduces to $\mathbf{g}_{mk}=\overline{\mathbf{g}}_{mk}+\check{\mathbf{g}}_{mk}$, the closed-form results in Theorem 1 and Theorem 3 would have different forms. More specifically, we obtain the expressions in Theorem 1, but using $\overline{\mathbf{G}}_{kk}=\overline{\mathbf{g}}_{k}^{\prime}\overline{\mathbf{g}}_{k}^{\prime,H}$ with $\overline{\mathbf{g}}_{k}^{\prime}\triangleq[\overline{\mathbf{g}}_{1k}^{T},\dots,\overline{\mathbf{g}}_{Mk}^{T}]^{T}$,
$\varepsilon_{kl}$ $=\sqrt{p_{k}p_{l}}\tau_{p}\mathrm{tr}(\mathbf{W}_{k}^{H}\overline{\mathbf{G}}_{lk})\mathrm{tr}(\mathbf{W}_{k}\check{\mathbf{R}}_{k}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{l})$
$+\sqrt{p_{k}p_{l}}\tau_{p}\mathrm{tr}(\mathbf{W}_{k}\overline{\mathbf{G}}_{kl})\mathrm{tr}(\mathbf{W}_{k}^{H}\check{\mathbf{R}}_{l}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{k})$
@@
with $\overline{\mathbf{G}}_{lk}=\overline{\mathbf{g}}_{l}^{\prime}\overline{\mathbf{g}}_{k}^{\prime,H}$, and $\omega_{k}=0$, respectively. Furthermore, the expressions in Theorem 3 hold but using $\mathbf{\Gamma}_{k}=\sum_{l=1}^{K}{p_{l}[(\overline{\mathbf{G}}_{kk}^{T}\otimes\overline{\mathbf{G}}_{ll})+(\overline{\mathbf{G}}_{kk}^{T}\otimes\check{\mathbf{R}}_{l})+(\widehat{\mathbf{R}}_{k}^{T}\otimes\overline{\mathbf{G}}_{ll})}+(\widehat{\mathbf{R}}_{k}^{T}\otimes\check{\mathbf{R}}_{l})]+\sum_{l\in\mathcal{P}_{k}}{p_{l}[\sqrt{p_{k}p_{l}}\tau_{p}\overline{\mathbf{g}}_{lk}\tilde{\mathbf{r}}_{lk}^{H}+\sqrt{p_{k}p_{l}}\tau_{p}\tilde{\mathbf{r}}_{lk}\overline{\mathbf{g}}_{lk}^{H}}+p_{k}p_{l}\tau_{p}^{2}\tilde{\mathbf{r}}_{lk}\tilde{\mathbf{r}}_{lk}^{H}]-p_{k}\overline{\mathbf{r}}_{k}\overline{\mathbf{r}}_{k}^{H}+\sigma^{2}(\overline{\mathbf{R}}_{k}^{T}\otimes\mathbf{I}_{MN})$, where $\overline{\mathbf{g}}_{lk}=\mathrm{vec}(\overline{\mathbf{G}}_{lk})$. These results can be easily derived following the same methods as in Appendix B and Appendix E. As observed, if the phase-shifts are perfectly mitigated, some cross terms from different APs or different UEs become non-zero, such as $\overline{\mathbf{G}}_{lk}$. Note that the results for the distributed processing scheme, including the achievable SE, DG-OBE, and DL-OBE expressions in scenarios with perfectly mitigated phase-shifts, can also be derived using the same approach as this remark. Therefore, they are omitted in the subsequent discussion for brevity.
@@ IV-B DG-OBE Combining Scheme Design
In this subsection, we study the DG-OBE combining scheme design to maximize (14). We define $\mathbf{w}_{k}^{\mathrm{d}}=[\mathbf{w}_{1k}^{T},\mathbf{w}_{2k}^{T},\dots,\mathbf{w}_{Mk}^{T}]^{T}\in\mathbb{C}^{MN^{2}}$, where $\mathbf{w}_{mk}=\mathrm{vec}(\mathbf{W}_{mk})\in\mathbb{C}^{N^{2}}$. To separate the optimization of $\{\mathbf{W}_{mk}:m=1,\dots,M\}$ and $\mathbf{a}_{k}$, we first assume all elements in $\mathbf{a}_{k}$ to be equal and thus, $\mathrm{SINR}_{k}^{\mathrm{LSFD}}$ in (15) reduces to (26) on the top of the previous page. Based on (26), we can optimize the distributed BE matrices $\{\mathbf{W}_{mk}:m=1,\dots,M\}$ to maximize the achievable SE performance for UE $k$ in (26) as follows.
@@
The optimal matrices $\{\mathbf{W}_{mk}^{*}:m=1,\dots,M\}$, maximizing the effective SINR for the distributed processing scheme in (26), can be computed as $\mathbf{W}_{mk}^{*}=\mathrm{vec}^{-1}(\mathbf{w}_{mk}^{\mathrm{d},*})\in\mathbb{C}^{N\times N}$ with $\mathbf{w}_{mk}^{\mathrm{d},*}=[\mathbf{w}_{k}^{\mathrm{d},*}]_{(m-1)N^{2}+1:mN^{2}}\in\mathbb{C}^{N^{2}}$ and
$\mathbf{w}_{k}^{\mathrm{d},*}=$ (27)
$\left(\sum_{l=1}^{K}{p_{l}\mathbb{E}\{\mathbf{z}_{kl}\mathbf{z}_{kl}^{H}\}-p_{k}\mathbb{E}\{\mathbf{h}_{k}\}\mathbb{E}\{\mathbf{h}_{k}\}^{H}+\sigma^{2}\mathbf{\Theta}_{k}}\right)^{-1}\mathbb{E}\{\mathbf{h}_{k}\},$
where $\mathbf{z}_{kl}=[\mathrm{vec}(\mathbf{g}_{1l}\widehat{\mathbf{g}}_{1k}^{H})^{T},\dots,\mathrm{vec}(\mathbf{g}_{Ml}\widehat{\mathbf{g}}_{Mk}^{H})^{T}]^{T}\in\mathbb{C}^{MN^{2}}$, $\mathbf{h}_{k}=[\mathrm{vec}(\mathbf{g}_{1k}\widehat{\mathbf{g}}_{1k}^{H})^{T},\dots,\mathrm{vec}(\mathbf{g}_{Mk}\widehat{\mathbf{g}}_{Mk}^{H})^{T}]^{T}\in\mathbb{C}^{MN^{2}}$, and $\mathbf{\Theta}_{k}=\mathrm{diag}\{(\mathbb{E}\{\widehat{\mathbf{g}}_{1k}\widehat{\mathbf{g}}_{1k}^{H}\}^{T}\otimes\mathbf{I}_{N}),\dots,(\mathbb{E}\{\widehat{\mathbf{g}}_{Mk}\widehat{\mathbf{g}}_{Mk}^{H}\}^{T}\otimes\mathbf{I}_{N})\}\in\mathbb{C}^{MN^{2}\times MN^{2}}$. When $\{\mathbf{W}_{mk}^{*}:m=1,\dots,M\}$ are applied, the SINR in (26) can achieve the maximum value as $\widetilde{\mathrm{SINR}}_{k}^{\mathrm{LSFD},*}=p_{k}\mathbb{E}\{\mathbf{h}_{k}\}^{H}\mathbf{w}_{k}^{\mathrm{d},*}$.
@@
The closed-form results of $\mathbf{w}_{k}^{\mathrm{d},*}$ can be computed as $\overline{\mathbf{w}}_{k}^{\mathrm{d},*}=\mathbf{\Lambda}_{k}^{-1}\overline{\mathbf{h}}_{k}$, where $\mathbf{\Lambda}_{k}=\sum_{l=1}^{K}{\mathbf{\Lambda}_{kl}^{(1)}}+\sum_{l\in\mathcal{P}_{k}}{(\mathbf{\Lambda}_{kl}^{(2)}+\mathbf{\Lambda}_{kl}^{(3)})}-p_{k}\overline{\mathbf{h}}_{k}\overline{\mathbf{h}}_{k}^{H}+\sigma^{2}\mathbf{\Lambda}_{k}^{(4)}\in\mathbb{C}^{MN^{2}\times MN^{2}}$ and $\overline{\mathbf{h}}_{k}=[\overline{\mathbf{h}}_{1k}^{T},\dots,\overline{\mathbf{h}}_{Mk}^{T}]^{T}\in\mathbb{C}^{MN^{2}}$ with $\overline{\mathbf{h}}_{mk}=\mathrm{vec}(\overline{\mathbf{R}}_{mk})\in\mathbb{C}^{N^{2}}$. Besides, $\mathbf{\Lambda}_{kl}^{(1)}=\mathrm{diag}(\mathbf{\Lambda}_{1kl}^{(1)},\dots,\mathbf{\Lambda}_{Mkl}^{(1)})$, $\mathbf{\Lambda}_{kl}^{(2)}=\mathrm{diag}(\mathbf{\Lambda}_{1kl}^{(2)},\dots,\mathbf{\Lambda}_{Mkl}^{(2)})$, $\mathbf{\Lambda}_{k}^{(4)}=\mathrm{diag}(\mathbf{\Lambda}_{1k}^{(4)},\dots,\mathbf{\Lambda}_{Mk}^{(4)})$ are all diagonal block matrices with dimension $\mathbb{C}^{MN^{2}\times MN^{2}}$, where $\mathbf{\Lambda}_{mkl}^{(1)}=p_{l}[(\overline{\mathbf{G}}_{mkk}^{T}\otimes\overline{\mathbf{G}}_{mll})+(\overline{\mathbf{G}}_{mkk}^{T}\otimes\check{\mathbf{R}}_{ml})+(\widehat{\mathbf{R}}_{mk}^{T}\otimes\overline{\mathbf{G}}_{mll})+(\widehat{\mathbf{R}}_{mk}^{T}\otimes\check{\mathbf{R}}_{ml})]$,